Please show work, I am trying to learn the best I can, thank you in advance!
4. An electron gun, operating in a vacuum, shoots electrons at a metal plate that is 4.0 mm away. The plate is 5.0 volts lower in potential than the gun. What velocity is required for the electons to reach the plate. thnt:aner should be behwzen 1000 and 1500
Answer #1

from work energy conservation theorem

W = KEi +KEf

The electrons accelerated through a potential difference of V , so work done is V*q.

V*q = 0 + 1/2*m*v2

v = sqrt (2Vq/m)

v = sqrt (2*5*1.6*10-19 / 9.1*10-31)

v = 1.32*106 m/s

so velocity required for the electron to reach the plate is 1.32*106 m/s.